# Kusatsu, Shiga 525-8577, Japan

### v1

@@ D.2 SCSMs with boundaries
${\{{{S},{S}}\}}$ $=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho_{a}^{i}{\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
@@
$=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+\frac{1}{2}m_{ab}{\bm{A}}^{a}\mathrm{d}{\bm{A}}^{b}+{\bm{Z}}_{i}{\bm{Q}}^{i}-\frac{1}{2}\omega^{c}_{ai}m_{cb}{\bm{A}}^{a}{\bm{A}}^{b}{\bm{Q}}^{i}\right$.
@@
$S_{b}$ $=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[\frac{1}{2}\mu^{(2)}_{ij}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}+\mu^{(1)}_{ia}(\bm{X}){\bm{Q}}^{i}{\bm{A}}^{a}+\frac{1}{2}\mu^{(0)}_{ab}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}\right].$ (243)
@@
$\delta\bm{X}^{i}$ $={\{{{S},{\bm{X}^{i}}}\}}=\mathrm{d}\bm{X}^{i}-{\bm{Q}}^{i}-\rho^{i}_{a}{\bm{A}}^{a},$ (244)
@@ E.1 Introduction of fluxes
$S_{F}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}H^{(2)}_{ABa}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{Aab}(\bm{X}){\bm{Y}}^{A}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (267)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\bm{W}}_{a}D{\bm{A}}^{a}\right$.
$\left.+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}+\frac{1}{3!}H_{ABC}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}+{\rho_{(A)}}^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}-\frac{1}{2}T_{ab}^{c}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right$.
$\left.+\frac{1}{2}H^{(2)}_{ABa}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{Aab}(\bm{X}){\bm{Y}}^{A}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (268)
@@ E.2 General CSMs with boundaries
${\{{{S},{S}}\}}$ $=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\bm{W}}_{a}D{\bm{A}}^{a}\right$.
@@
$=2\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\bm{W}}_{a}\mathrm{d}{\bm{A}}^{a}+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}{\bm{Y}}^{A}\right$.
@@
$S_{b}$ $=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[\frac{1}{2}\mu^{(2)}_{AB}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}+\mu^{(1)}_{Aa}(\bm{X}){\bm{Y}}^{A}{\bm{A}}^{a}+\frac{1}{2}\mu^{(0)}_{ab}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}\right],$ (270)
where $\mu^{(2)}$ is an $E$ two-form, $\mu^{(1)}$ is an $E$ one-form taking a values on $A$, $\mu^{(2)}$ is an $E$ zero-form taking a values on $\wedge^{2}A$.
@@
$\delta\bm{X}^{i}$ $={\{{{S},{\bm{X}^{i}}}\}}=\mathrm{d}\bm{X}^{i}-{\rho_{(E)}}^{i}_{A}{\bm{Y}}^{A}-{\rho_{(A)}}^{i}_{a}{\bm{A}}^{a},$ (271)
@@ F.1 Introduction of fluxes
$S_{F}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}H^{(2)}_{ABa}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{Aab}(\bm{X}){\bm{Y}}^{A}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (297)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}(\bm{X}){\bm{A}}^{a}D{\bm{A}}^{b}\right$.
$\left.+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}+\frac{1}{3!}H_{ABC}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}+{\rho_{(A)}}^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}+\frac{1}{3!}T_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right$.
$\left.+\frac{1}{2}H^{(2)}_{ABa}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{Aab}(\bm{X}){\bm{Y}}^{A}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (298)
@@ F.2 General CSMs with boundaries
${\{{{S},{S}}\}}$ $=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}\right$.
@@
$=2\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\bm{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}{\bm{A}}^{a}\mathrm{d}{\bm{A}}^{b}\right$.
@@
$S_{b}$ $=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[\frac{1}{2}\mu^{(2)}_{AB}(\bm{X}){\bm{Y}}^{A}{\bm{Y}}^{B}+\mu^{(1)}_{Aa}(\bm{X}){\bm{Y}}^{A}{\bm{A}}^{a}+\frac{1}{2}\mu^{(0)}_{ab}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}\right].$ (300)
@@
$\delta\bm{X}^{i}$ $={\{{{S},{\bm{X}^{i}}}\}}=\mathrm{d}\bm{X}^{i}-{\rho_{(E)}}^{i}_{A}{\bm{Y}}^{A}-{\rho_{(A)}}^{i}_{a}{\bm{A}}^{a},$ (301)

### v2

@@ D.2 SCSMs with boundaries
${\{{{S},{S}}\}}$ $=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho_{a}^{i}{\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
@@
$=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+\frac{1}{2}m_{ab}{\bm{A}}^{a}\mathrm{d}{\bm{A}}^{b}+{\bm{Z}}_{i}{\bm{Q}}^{i}-\frac{1}{2}\omega^{c}_{ai}m_{cb}{\bm{A}}^{a}{\bm{A}}^{b}{\bm{Q}}^{i}\right$.
@@
$S_{b}$ $=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[\frac{1}{2}\mu^{(2)}_{ij}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}+\mu^{(1)}_{ia}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{A}}^{a}+\frac{1}{2}\mu^{(0)}_{ab}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}\right].$ (243)
@@
$\delta\boldsymbol{X}^{i}$ $={\{{{S},{\boldsymbol{X}^{i}}}\}}=\mathrm{d}\boldsymbol{X}^{i}-{\bm{Q}}^{i}-\rho^{i}_{a}{\bm{A}}^{a},$ (244)
@@ E.1 Introduction of fluxes
$S_{F}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}H^{(2)}_{ABa}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{Aab}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (267)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\bm{W}}_{a}D{\bm{A}}^{a}\right$.
$\left.+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}+\frac{1}{3!}H_{ABC}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}+{\rho_{(A)}}^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}-\frac{1}{2}T_{ab}^{c}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right$.
$\left.+\frac{1}{2}H^{(2)}_{ABa}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{Aab}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (268)
@@ E.2 General CSMs with boundaries
${\{{{S},{S}}\}}$ $=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\bm{W}}_{a}D{\bm{A}}^{a}\right$.
@@
$=2\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+{\bm{W}}_{a}\mathrm{d}{\bm{A}}^{a}+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}{\bm{Y}}^{A}\right$.
@@
$S_{b}$ $=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[\frac{1}{2}\mu^{(2)}_{AB}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}+\mu^{(1)}_{Aa}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{A}}^{a}+\frac{1}{2}\mu^{(0)}_{ab}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}\right],$ (270)
where $\mu^{(2)}$ is an $E$ two-form, $\mu^{(1)}$ is an $E$ one-form taking a values on $A^{*}$, $\mu^{(2)}$ is an $E$ zero-form taking a values on $\wedge^{2}A^{*}$.
@@
$\delta\boldsymbol{X}^{i}$ $={\{{{S},{\boldsymbol{X}^{i}}}\}}=\mathrm{d}\boldsymbol{X}^{i}-{\rho_{(E)}}^{i}_{A}{\bm{Y}}^{A}-{\rho_{(A)}}^{i}_{a}{\bm{A}}^{a},$ (271)
@@ F.1 Introduction of fluxes
$S_{F}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}H^{(2)}_{ABa}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{Aab}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (297)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}(\boldsymbol{X}){\bm{A}}^{a}D{\bm{A}}^{b}\right$.
$\left.+{\rho_{(E)}}^{i}_{A}{\bm{Z}}_{i}^{\nabla}{\bm{Y}}^{A}+\frac{1}{3!}H_{ABC}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{Y}}^{C}+{\rho_{(A)}}^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}+\frac{1}{3!}T_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right$.
$\left.+\frac{1}{2}H^{(2)}_{ABa}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{Aab}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (298)
@@ F.2 General CSMs with boundaries
${\{{{S},{S}}\}}$ $=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}\right$.
@@
$=2\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\boldsymbol{X}^{i}+\frac{1}{2}k_{AB}{\bm{Y}}^{A}\mathrm{d}{\bm{Y}}^{B}+\frac{1}{2}m_{ab}{\bm{A}}^{a}\mathrm{d}{\bm{A}}^{b}\right$.
@@
$S_{b}$ $=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[\frac{1}{2}\mu^{(2)}_{AB}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{Y}}^{B}+\mu^{(1)}_{Aa}(\boldsymbol{X}){\bm{Y}}^{A}{\bm{A}}^{a}+\frac{1}{2}\mu^{(0)}_{ab}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}\right].$ (300)
@@
$\delta\boldsymbol{X}^{i}$ $={\{{{S},{\boldsymbol{X}^{i}}}\}}=\mathrm{d}\boldsymbol{X}^{i}-{\rho_{(E)}}^{i}_{A}{\bm{Y}}^{A}-{\rho_{(A)}}^{i}_{a}{\bm{A}}^{a},$ (301)